Jackie is wrapping a birthday present. The gift box is in the shape of a rectangular prism. It has a length of 8 inches, a width of 10 inches, and a height of 3 inches. What is the minimum amount of wrapping paper she will need to completely cover the box?
Answer by ikleyn(52787)

.
Jackie is wrapping a birthday present. The gift box is in the shape of a rectangular prism.
It has a length of 8 inches, a width of 10 inches, and a height of 3 inches. What is the minimum amount of
wrapping paper she will need to completely cover the box?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

If the condition means the surface area of the box, then it is 2*(8*10 + 8*3 + 10*3) = . . .

The surface area of the box with dimensions AxBxC linear units is 2*(AB + AC + BC) square units.

If the condition means conventional/regular rectangular piece of paper, then it is, probably,
rectangle of 8+3+3 = 14 inches wide and 3+10+3+10 = 26 inches long.


Note: I used to think that the length is always more than the width.
It is not so in your condition.
It is not fatal, but creates unnecessary questions.